in going from one city to another, a driver tend to get lost travel 30 kms north,then 50 kms. west,20 kms. southeast. what is the distance between the two cities?
Answer by rfer(16322)

c^2=50^2+30^2
c^2=3400
c=58.3
58.3-20=38.3